# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 6.1 Klein four group case
First, we assume that the hyperfocal subalgebra $\mathbb{A}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$. Then the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}M$ and its dual induce a Morita equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$, where $M=\Omega_{\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}}(\mathcal{O}L)^{\circ}}^{n}(\mathbb{A}_{\delta})$. By Lemma 5.18, there is an element $a_{u,\delta}$ of $A_{\delta}^{*}$ such that it acts trivially on the subalgebra $\mbox{$\mathcal{O}$}L$ of $\mathbb{A}_{\delta}$ through the conjugation action on $\mathbb{A}_{\delta}$. For any $\mathbb{A}_{\delta}$-$\mbox{$\mathcal{O}$}L$-bimodule $U$, we can define the twisted bimodule ${{}_{(u,u)}}U$ by setting ${{}_{(u,u)}}U=U$ as an $\mathcal{O}$-module, with $a_{u,\delta}\mathfrak{a}a_{u,\delta}^{-1}\otimes_{\mathcal{O}}y$ acting on $\mathrm{u}\in{{}_{(u,u)}}U$ as $\mathfrak{a}\otimes_{\mathcal{O}}y$ on $\mathrm{u}\in U$, where $\mathfrak{a}$ belongs to $\mathbb{A}_{\delta}$ and $y$ belongs to $L$. Then it is easy to check that its conjugation action on $\mathbb{A}_{\delta}$ can induce an $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule isomorphism between $\mathbb{A}_{\delta}$ and ${{}_{(u,u)}}\mathbb{A}_{\delta}$. By the construction of the Heller operator $\Omega$, we can get an $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule isomorphism $\vartheta:M\longrightarrow{{}_{(u,u)}}M$. Now we can define a map
@@
$\varTheta:\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}M\longrightarrow{{}_{(u,u)}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}M),~~~~aj\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}m\mapsto uaja_{u,\delta}^{-1}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}\vartheta(m),$
@@
for any $aj\in\mathbb{A}j$ and any $m\in M$. Then it is routine to check that this map is well-defined and gives an $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule isomorphism. In particular, this Morita equivalence is $\Delta(P_{0})$-equivariant.
@@
Consier the second case where $\mathbb{A}$ is Morita equivalent to the principal block $B_{0}(\mathrm{A}_{5})$ of $\mathrm{A}_{5}$. We have an analogue notation ${{}_{(u,u)}}U^{\prime}$ for any $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule $U^{\prime}$. By [16, Theorem 4.14.2], the $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule $N^{\prime}$, which is an indecomposable direct summand of $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}L}B_{0}(\mathrm{A}_{5})$, is unique up to isomorphisms. It is clear that $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}L}B_{0}(\mathrm{A}_{5})$ is isomorphic to $){{}_{(u,u)}}(\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}L}B_{0}(\mathrm{A}_{5}$)) as $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodules. Therefore, we have an $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule isomorphism $\vartheta^{\prime}:N\longrightarrow{{}_{(u,u)}}N$ with the same argument above. Similar to the construction of map $\varTheta$, we can get an $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule isomorphism $\varTheta^{\prime}:\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}N\longrightarrow{{}_{(u,u)}}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}N$, sending $aj\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}n$ to $uaja_{u,\delta}^{-1}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}\vartheta^{\prime}(n)$ for any $aj\in\mathbb{A}j$ and any $n\in N$. We are done. ∎
Corollary 6.12.
The actions of $P_{0}$ on $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$ and $\mathrm{Irr}_{k}(\mathbb{A})$ are trivial.
Proof.
This follows from Lemma 6.11 easily. ∎
6.13.
@@
6.14.
There are two possibilities on $|\mathrm{Irr}_{k}(A)|$ (see [7, Theorem 1.1]). Let us recall some notation and results from [7]. Suppose that $P_{0}=P$, namely, $D$ is in the center of $P$. Then $N_{G}(P_{\gamma})$ controls fusion of the block $b$ and $|\mathrm{Irr}_{k}(A)|=3=|\mathrm{Irr}_{k}(\mathbb{A})|$. Furthermore, by the proof of Corollary 5.21, we can identify $E$ with the inertial quotient $N_{G}(P_{\gamma})/PC_{G}(P)$ and then $P=D\times C_{P}(E)$. We denote $C_{P}(E)$ by $R$ in this case. The other case is that $P_{0}<P$. In this case, $|\mathrm{Irr}_{k}(A)|=2$ and $|P:P_{0}|=2$. Furthermore, $E$ can be identified with a subgroup of $\mathrm{Aut}(P_{0})$ and then $P_{0}=D\times C_{P_{0}}(E)$ with this identification. We denote $C_{P_{0}}(E)$ by $R_{0}$ in this case.
Now we have the following structures about $\mathrm{Irr}_{\mathcal{K}}(A)$ and $\mathrm{Irr}_{k}(A)$.
@@
(ii) Suppose that $P_{0}<P$. Then we can choose notation such that $\bar{W}_{1}$ is $P$-stable and the stabilizer of $\bar{W}_{2}$ in $P$ is $P_{0}$. So $P$ can permute $\bar{W}_{2}$ and $\bar{W}_{3}$. Moreover, $\bar{W}_{1}$ can be extended to $\bar{A}$ and $\bar{W}_{\mathrm{i}}$ can be extended to $\bar{\mathbb{A}}_{P_{0}}$ for $\mathrm{i}=1,2$. For $\mathrm{Irr}_{\mathcal{K}}(A)$, we can obtain that $\hat{U}_{1}$ and $\hat{U}_{4}$ are $P$-stable and the stabilizer of $\hat{U}_{2}$ in $P$ is $P_{0}$. So $P$ can permute $\hat{U}_{2}$ and $\hat{U}_{3}$. Moreover, $\hat{U}_{1}$ and $\hat{U}_{4}$ can be extended to $\hat{A}$ and $\hat{U}_{\mathrm{i}}$ can be extended to $\hat{\mathbb{A}}_{P_{0}}$ for $\mathrm{i}=1,2$. In conclusion, we have
@@
For the case (i), since $|\mathrm{Irr}_{k}(A)|=3=|\mathrm{Irr}_{k}(\mathbb{A})|$ and $P$ is a $2$-group, it is easy to see that every simple $\bar{\mathbb{A}}$-module is $P$-stable. Then its extendibility is well known since $P$ is a $2$-group. The description of $\mathrm{Irr}_{k}(A)$ can be deduced from this extendibility. Consider simple $\hat{A}-$modules. Obviously, the decomposition map $d_{\mathbb{A}}$ commutes with the action of $P$. Then by equations (6.10.1) and (6.10.2), we can easily obtain that every simple $\hat{\mathbb{A}}-$module is $P$-stable. By Proposition 6.9, $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mathrm{i}})$ is coprime to $2$ for any $\mathrm{i}\in\{1,2,3,4\}$. Then by [13, Proposition 2.4], every simple $\hat{\mathbb{A}}-$module can be extended to $\hat{A}$. Hence, the description of $\mathrm{Irr}_{\mathcal{K}}(A)$ follows from Clifford Theory and the fact that $P/D$ is isomorphic to $R$.
For the case (ii), by Corollary 6.12, every simple $\hat{\mathbb{A}}-$module and every simple $\bar{\mathbb{A}}$-module are both $P_{0}$-stable. So the actions of $P$ on these two sets are essentially equivalent to the actions of $P/P_{0}$, which has order $2$. Since in this case $|\mathrm{Irr}_{k}(A)|=2$ and $|\mathrm{Irr}_{k}(\mathbb{A})|=3$, then one of simple $\bar{\mathbb{A}}$-modules is $P$-stable and $P/P_{0}$ permutes the remaining two simple modules. When $\mathbb{A}$ is Morita equivalent to the principal block $B_{0}(\mathrm{A}_{5})$ of $\mathrm{A}_{5}$, by the equations (6.10.2), $\bar{W}_{1}$ has to be the unique $P$-stable simple $\bar{\mathbb{A}}$-module. When $\mathbb{A}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$, we have to choose notation such that $\bar{W}_{1}$ is $P$-stable since we do not known whether the Morita equivalence in this case is $\Delta{P}$-equivariant. Then $P/P_{0}$ permutes $\bar{W}_{2}$ and $\bar{W}_{3}$. Similarly, we can obtain that $\bar{W}_{1}$ can be extended to $\bar{A}$ and $\bar{W}_{\mathrm{i}}$ can be extended to $\bar{\mathbb{A}}_{P_{0}}$ for $\mathrm{i}=2,3$. Now the description of $\mathrm{Irr}_{k}(A)$ follows from the extendibility and Clifford Theory. Consider simple $\hat{A}-$modules. By equations (6.10.1) and (6.10.2), we can obtain that $\hat{U}_{1}$ and $\hat{U}_{4}$ are $P$-stable and $P/P_{0}$ can permute $\hat{U}_{2}$ and $\hat{U}_{3}$. With the same argument above, $\hat{U}_{1}$ and $\hat{U}_{4}$ can be both extended to $\hat{A}$. Note that the proof of [13, Proposition 2.4] is still valid if we replace $P$ and $\hat{A}$ there by $P_{0}$ and $\hat{\mathbb{A}}_{P_{0}}$. Then $\hat{U}_{2}$ and $\hat{U}_{3}$ can be both extended to $\hat{\mathbb{A}}_{P_{0}}$. Similarly, the description of $\mathrm{Irr}_{\mathcal{K}}(A)$ follows from Clifford Theory and the fact that $P_{0}/D$ is isomorphic to $R_{0}$. ∎
@@ 6.2 Cyclic group case
In this subsection, we will assume the hyperfocal subgroup $D$ is a nontrivial cyclic group. In this case the hyperfocal quotient inertial $E_{\mathfrak{h}}$ is a cyclic group of order divides p1p-1 and acts freely on $D-\{1\}$. Then we can apply Theorem 5.14 to this case and get a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. Here $L=D\rtimes E_{\mathfrak{h}}$. It is well-known that this stable equivalence of Morita type can be lifted to a Rickard equivalence by adopting the method in the proof of [17, Theorem 11.12.1]. In the following, we will investigate the equivariant property of this Rickard equivalence.
6.16.
@@
Since the hyperfocal subgroup $D$ is cyclic, by [33, Theorem 3], $N_{G}(P_{\gamma})$ controls fusion of the block $b$. So we can identify $E_{\mathfrak{h}}$ with the inertial quotient $E=N_{G}(P_{\gamma})/PC_{G}(P)$ and then $P=D\rtimes C_{P}(E)$. We denote $C_{P}(E)$ by $R$. Then the conjutation action of $R$ on $E$ is trivial. Moreover, by [33, Theorem 1], $l(b)=|\mathrm{IBr}(b)|=|E|$. Then $l(A)$ also equals $|E|$. On the other hand, since $A$ is symmetric and stably equivalent to $\mbox{$\mathcal{O}$}L$, which is a serial algebra in this case, by [17, Proposition 11.6.1], $l(\mathbb{A})=|\mathrm{IBr}(L)|$, which is just $|E|$. Hence, we have $l(\mathbb{A})=l(A)$ equal to the inertial index $|E|$ of the block. It can be deduced from this equality that $P$ acts trivially on $\mathrm{Irr}_{k}(\mathbb{A})$ and then every simple $\bar{\mathbb{A}}$-module can be extended to $\bar{A}$.
6.17.
@@
Let $\mathfrak{i}$ and $\varsigma$ be primitive idempotents of $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$, respectively. For any $u\in P$, we have the twisted $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule ${{}_{(u,u)}}(\mathbb{A}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma\mbox{$\mathcal{O}$}L)$. It is clear that the map sending $\mathfrak{a}\mathfrak{i}^{u}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma^{u}x$ to ${{}^{u}}\mathfrak{a}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma{{}^{u}}x$ defines an isomorphism of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules from $\mathbb{A}\mathfrak{i}^{u}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma^{u}\mbox{$\mathcal{O}$}L$ to ${{}_{(u,u)}}(\mathbb{A}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma\mbox{$\mathcal{O}$}L)$, where $\mathfrak{a}\in\mathbb{A}$ and $x\in L$. By the paragraph above, $P$ acts trivially on $\mathrm{Irr}_{k}(\mathbb{A})$. It is clear that $\mathbb{A}\mathfrak{i}^{u}$ is isomorphic to $\mathbb{A}\mathfrak{i}$ as left $\mathbb{A}$-modules. Similarly, $P$ acts on $\mathrm{IBr}(L)$ trivially and then $\varsigma^{u}\mbox{$\mathcal{O}$}L$ is isomorphic to $\varsigma\mbox{$\mathcal{O}$}L$ as right $\mbox{$\mathcal{O}$}L$-modules. In conclusion, ${{}_{(u,u)}}(\mathbb{A}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma\mbox{$\mathcal{O}$}L)$ is isomorphic to $\mathbb{A}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma\mbox{$\mathcal{O}$}L$ for any $u\in P$. Hence, we have the following equivariant property for projective $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules.
@@
6.19.
Let $M_{\cdot}=(M_{n},\partial_{n})$ be a bounded complex of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. For any $u\in P$, we can define the twisted complex ${{}_{(u,u)}}(M_{\cdot})$ with ${{}_{(u,u)}}(M_{\cdot})_{n}={{}_{(u,u)}}(M_{n})$ and the morphism from ${{}_{(u,u)}}(M_{\cdot})_{n}$ to ${{}_{(u,u)}}(M_{\cdot})_{n-1}$ is just $\partial_{n}$. The complex $M_{\cdot}$ is called $P$-invariant if $M_{\cdot}$ is isomorphic to ${{}_{(u,u)}}(M_{\cdot})$ in $\mathcal{D}^{b}(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})$. Furthermore, if the complex $M_{\cdot}$ induces a derived equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$, then this derived equivalence is called $P$-equivariant. We refer to [18, Definition 2.5] for these two definitions.
@@
There is a $P$-invariant Rickard $2$-term complex $\mathbb{M}_{\cdot}$ of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. In particular, there is a $P$-equivariant Rickard equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$.

### v2

@@ 6.1 Klein four group case
First, we assume that the hyperfocal subalgebra $\mathbb{A}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$. Then the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}M$ and its dual induce a Morita equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$, where $M=\Omega_{\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}}(\mathcal{O}L)^{\circ}}^{n}(\mathbb{A}_{\delta})$. Then it is easy to check that the conjugation action of $a_{u,\delta}$ on $\mathbb{A}_{\delta}$ can induce an $\mathbb{A}_{\delta}$-$\mbox{$\mathcal{O}$}L$-bimodule isomorphism between $\mathbb{A}_{\delta}$ and ${{}_{(u,u)}}\mathbb{A}_{\delta}$. By the construction of the Heller operator $\Omega$, we can get an $\mathbb{A}_{\delta}$-$\mbox{$\mathcal{O}$}L$-bimodule isomorphism $\vartheta:M\longrightarrow{{}_{(u,u)}}M$. Now we can define a map
@@
$\varTheta:\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}M\longrightarrow{{}_{(u,u)}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}M),~~~~aj\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}m\mapsto u^{-1}aja_{u,\delta}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}\vartheta(m),$
@@
for any $aj\in\mathbb{A}j$ and any $m\in M$. Then it is routine to check that this map is well-defined and gives an $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule isomorphism. In particular, this Morita equivalence is $\Delta(P)$-equivariant.
@@
The remaining case is where $\mathbb{A}$ is Morita equivalent to the principal block $B_{0}(\mathrm{A}_{5})$ of $\mathrm{A}_{5}$. We have the analogue notation ${{}_{(u,u)}}U^{\prime}$ and ${{}_{(u,u)}}M^{\prime}$ for any $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule $U^{\prime}$ and any $\mathbb{A}$-$B_{0}(\mathrm{A}_{5})$-bimodule $M^{\prime}$, respectively. By [15, Theorem 4.14.2], the $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule $N^{\prime}$, which is an indecomposable direct summand of $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}L}B_{0}(\mathrm{A}_{5})$, is unique up to isomorphism. Similarly, the conjugation actions of $a_{u,\delta}$ on $\mathbb{A}_{\delta}$ and $u$ on $B_{0}({\rm A}_{5})$ can induce an isomorphism between $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}L}B_{0}(\mathrm{A}_{5})$ and $){{}_{(u,u)}}(\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}L}B_{0}(\mathrm{A}_{5}$)) as $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodules. Therefore, by the uniqueness of $N^{\prime}$, we have an $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule isomorphism $\vartheta^{\prime}:N\longrightarrow{{}_{(u,u)}}N$ with the same argument above since $N=\Omega_{\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}}B_{0}(\mathrm{A}_{5})^{\circ}}^{n}(N^{\prime})$. Similar to the construction of map $\varTheta$, we can get an $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule isomorphism $\varTheta^{\prime}:\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}N\longrightarrow{{}_{(u,u)}}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}N$, sending $aj\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}n$ to $u^{-1}aja_{u,\delta}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}\vartheta^{\prime}(n)$ for any $aj\in\mathbb{A}j$ and any $n\in N$. We are done. ∎
6.15.
@@
6.16.
There are two possibilities on $|l(A)|$ (see [7, Theorem 1.1]). Let us recall some notation and results from [7]. Suppose that $P_{0}=P$, namely, $D$ is in the center of $P$. Then $N_{G}(P_{\gamma})$ controls fusion of the block $b$ and $|l(A)|=3=|l(\mathbb{A})|$. Furthermore, by the proof of [7, Lemma 2.4], we can identify $E_{\mathfrak{h}}$ with the inertial quotient $N_{G}(P_{\gamma})/PC_{G}(P)$ through the restriction and then $P=D\times C_{P}(E_{\mathfrak{h}})$. We denote $C_{P}(E_{\mathfrak{h}})$ by $R$ in this case. The other case is that $P_{0}<P$. In this case, $|l(A)|=2$ and $|P:P_{0}|=2$. Furthermore, $E_{\mathfrak{h}}$ can be identified with a subgroup of $\mathrm{Aut}(P_{0})$ and then $P_{0}=D\times C_{P_{0}}(E_{\mathfrak{h}})$ with this identification. We denote $C_{P_{0}}(E_{\mathfrak{h}})$ by $R_{0}$ in this case.
Now we can get the structures about $\mathrm{Irr}_{\mathcal{K}}(A)$ and $\mathrm{Irr}_{k}(A)$ as follows.
@@
(ii) Suppose that $P_{0}<P$. Then $\bar{W}_{1}$ is $P$-stable and the stabilizer of $\bar{W}_{2}$ in $P$ is $P_{0}$. So the action of $P$ can permute $\bar{W}_{2}$ and $\bar{W}_{3}$. Moreover, $\bar{W}_{1}$ can be extended to $\bar{A}$ and $\bar{W}_{\mathrm{i}}$ can be extended to $\bar{\mathbb{A}}_{P_{0}}$ for $\mathrm{i}=1,2$. For $\mathrm{Irr}_{\mathcal{K}}(A)$, we can obtain that $\hat{U}_{1}$ and $\hat{U}_{4}$ are $P$-stable and the stabilizer of $\hat{U}_{2}$ in $P$ is $P_{0}$. So the action of $P$ can permute $\hat{U}_{2}$ and $\hat{U}_{3}$. Moreover, $\hat{U}_{1}$ and $\hat{U}_{4}$ can be extended to $\hat{A}$ and $\hat{U}_{\mathrm{i}}$ can be extended to $\hat{\mathbb{A}}_{P_{0}}$ for $\mathrm{i}=1,2$. In conclusion, we have
@@
For the case (i), by Proposition 6.14, it is easy to see that every simple $\bar{\mathbb{A}}$-module is $P$-stable. Then the extendibility is well known since $P$ is a $2$-group. The description of $\mathrm{Irr}_{k}(A)$ can be deduced from this extendibility. Let us consider simple $\hat{A}-$modules. Obviously, the decomposition map $d_{\mathbb{A}}$ commutes with the action of $P$. Then by the equations (6.12.1) and (6.12.2), we can easily obtain that every simple $\hat{\mathbb{A}}-$module is $P$-stable. By Proposition 6.11, $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mathrm{i}})$ is coprime to $2$ for any $\mathrm{i}\in\{1,2,3,4\}$. Then by [12, Proposition 2.4], every simple $\hat{\mathbb{A}}-$module can be extended to $\hat{A}$. Hence, the description of $\mathrm{Irr}_{\mathcal{K}}(A)$ follows from Clifford Theory and the fact that $P/D$ is isomorphic to $R$.
For the case (ii), by Proposition 6.14, every simple $\hat{\mathbb{A}}-$module and every simple $\bar{\mathbb{A}}$-module are both $P_{0}$-stable. Moreover, the action of $P$ can permute the two simple $\bar{\mathbb{A}}$-module $\bar{W}_{2}$ and $\bar{W}_{3}$. Then the description of the action of $P$ on ${\rm Irr}_{\mathcal{K}}(\mathbb{A})$ can be similarly obtained by the equations (6.12.1) and (6.12.2). With the same argument above, $\hat{U}_{1}$ and $\hat{U}_{4}$ can be both extended to $\hat{A}$. Note that the proof of [12, Proposition 2.4] is still valid if we replace $P$ and $\hat{A}$ there by $P_{0}$ and $\hat{\mathbb{A}}_{P_{0}}$. Then $\hat{U}_{2}$ and $\hat{U}_{3}$ can be both extended to $\hat{\mathbb{A}}_{P_{0}}$. Similarly, the description of $\mathrm{Irr}_{\mathcal{K}}(A)$ follows from Clifford Theory and the fact that $P_{0}/D$ is isomorphic to $R_{0}$. ∎
@@ 6.2 Cyclic group case
In this subsection, we will assume the hyperfocal subgroup $D$ is a nontrivial cyclic group. In this case the hyperfocal quotient inertial $E_{\mathfrak{h}}$ is a nontrivial cyclic group of order divides p1p-1 and acts freely on $D-\{1\}$. Then we can apply Theorem 5.14 to this case and get a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. It is well-known that this stable equivalence of Morita type can be lifted to a Rickard equivalence by adopting the method in the proof of [16, Theorem 11.12.1]. In the following, we will investigate the equivariant property of this Rickard equivalence.
6.18.
@@
Since the hyperfocal subgroup $D$ is cyclic, by [33, Theorem 3], $N_{G}(P_{\gamma})$ controls fusion of the block $b$. So we can identify $E_{\mathfrak{h}}$ with the inertial quotient $E=N_{G}(P_{\gamma})/PC_{G}(P)$ and then $P=D\rtimes C_{P}(E)$. We denote $C_{P}(E)$ by $R$. Then the conjutation action of $R$ on $E$ is trivial. Moreover, by [33, Theorem 1], $l(b)=|\mathrm{IBr}(b)|=|E|$. Then $l(A)$ also equals $|E|$. On the other hand, since $k\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$ is symmetric and stably equivalent to $kL$, which is a serial algebra in this case, by [16, Proposition 11.6.1], we have $l(\mathbb{A})=|\mathrm{IBr}(L)|=|E|$. Hence, we have $l(\mathbb{A})=l(A)$ equal to the inertial index $|E|$ of the block. It can be deduced from this equality that $P$ acts trivially on $\mathrm{Irr}_{k}(\mathbb{A})$ and then every simple $\bar{\mathbb{A}}$-module can be extended to $\bar{A}$.
6.19.
@@
Let $\mathfrak{i}$ and $\varsigma$ be primitive idempotents of $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$, respectively. For any $u\in P$, we have the twisted $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule ${{}_{(u,u)}}(\mathbb{A}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma\mbox{$\mathcal{O}$}L)$ (see the paragraph 4.9). It is clear that the map sending $\mathfrak{a}\mathfrak{i}^{u^{-1}}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma^{u^{-1}}x$ to $(\mathfrak{a}^{u})\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma x^{u}$ defines an isomorphism of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules from $\mathbb{A}\mathfrak{i}^{u}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma^{u}\mbox{$\mathcal{O}$}L$ to ${{}_{(u,u)}}(\mathbb{A}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma\mbox{$\mathcal{O}$}L)$, where $\mathfrak{a}\in\mathbb{A}$ and $x\in L$. By the paragraph above, $P$ acts trivially on $\mathrm{Irr}_{k}(\mathbb{A})$. It is clear that $\mathbb{A}\mathfrak{i}^{u^{-1}}$ is isomorphic to $\mathbb{A}\mathfrak{i}$ as left $\mathbb{A}$-modules. Similarly, $P$ acts on $\mathrm{IBr}(L)$ trivially and then $\varsigma^{u^{-1}}\mbox{$\mathcal{O}$}L$ is isomorphic to $\varsigma\mbox{$\mathcal{O}$}L$ as right $\mbox{$\mathcal{O}$}L$-modules. In conclusion, ${{}_{(u,u)}}(\mathbb{A}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma\mbox{$\mathcal{O}$}L)$ is isomorphic to $\mathbb{A}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma\mbox{$\mathcal{O}$}L$ for any $u\in P$. Hence, we have the following equivariant property for projective $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules.
@@
6.21.
Let $M_{\cdot}=(M_{n},\partial_{n})$ be a bounded complex of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. For any $u\in P$, we can define the twisted complex ${{}_{(u,u)}}(M_{\cdot})$ with $({{}_{(u,u)}}(M_{\cdot}))_{n}={{}_{(u,u)}}(M_{n})$ and the morphism from ${{}_{(u,u)}}(M_{\cdot})_{n}$ to ${{}_{(u,u)}}(M_{\cdot})_{n-1}$ is just $\partial_{n}$. The complex $M_{\cdot}$ is called $P$-invariant if $M_{\cdot}$ is isomorphic to ${{}_{(u,u)}}(M_{\cdot})$ in $\mathcal{D}^{b}(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})$. Furthermore, if the complex $M_{\cdot}$ induces a derived equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$, then this derived equivalence is called $P$-equivariant. We refer to [18, Definition 2.5] for these two definitions.
@@
There is a $P$-invariant 22-term Rickard complex $\mathbb{M}_{\cdot}$ of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. In particular, there is a $P$-equivariant Rickard equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$.